OpenStudy (anonymous):

amal is a safety inspector for a family gym. Recently, guests have been complaining about the diving board saying that it is leaning towards the water. Jamal needs to know whether this diving board is standing up straight or not. Jamal measures the height of the diving board to be 12 feet. The pool is 15 feet across and he measures the diagonal from the other side of the pool to the top of the diving board to be 19 feet. Is the diving board at a 90 degree angle with the ground? Explain how you know. Pythagoras theorem: a^2+b^2=c^2 12^2 + 15^2 ≠ 19^2 Is this accurate?

OpenStudy (anonymous):

This was my answer: Pythagoras theorem: a^2+b^2=c^2 12^2 + 15^2 ≠ 19^2 Is this accurate?

OpenStudy (callisto):

You haven't drawn the conclusion... :|

OpenStudy (anonymous):

well i was saying that you should try trigonometry i think it is not 90 if it does not follow the Pythagoras

OpenStudy (anonymous):

If: \[12^2 + 15^2 \ne 19^2\] That means angle is not 90.. And actually it is not 90 because they are not equal..

OpenStudy (callisto):

Using Pythagoras is fine ..... Becca has actually done the explain part (almost done) but not answering the question. Since 12^2 + 15^2 = 369 ≠ 19^2 Therefore, it is not an right-angled triangle (Converse of Pyth. Thm)

OpenStudy (anonymous):

Sorry guys, I was talking to my mom about something. Um, so did I do it right, or? I'm a little confused...

OpenStudy (anonymous):

http://www.teacherschoice.com.au/maths_library/trigonometry/solve_trig_sss.htm try this link if you want to know the exact angles

OpenStudy (anonymous):

I don't think I need to know the exact angles. I think I just need to know if it's 90º and how I know that.

OpenStudy (callisto):

Becca!~ Do you think it is a right angle? If you don't tell us, we can't check it for you!

OpenStudy (anonymous):

I don't think it is. Judging from Pythagoras Theorem they are not equal.

OpenStudy (callisto):

Then you're right :)

OpenStudy (anonymous):

:'3 Thanks Calli :D

OpenStudy (callisto):

Welcome!~